# Relative monodromy of abelian logarithms for finite covers of universal families

### v2

@@ 2.1 Setting
Let $S$ be a regular, irreducible, quasi-projective variety defined over $\mathbb{C}$. Let $\phi:\mathcal{A}\to S$ be an abelian scheme over $S$ of relative dimension $g\geq 1$, namely a proper smooth group $S$-scheme admitting a “zero section” $\sigma_{0}:S\to\mathcal{A}$ and whose fibers are $g$ dimensional abelian varieties. For any $s\in S(\mathbb{C})$ the fiber over $s$ is denoted by $\mathcal{A}_{s}$. We denote with $\Sigma^{\operatorname{an}}$ the sheaf over $S(\mathbb{C})$ of complex analytic sections of $\phi$ whereas $\Sigma(S)\subset\Sigma^{\operatorname{an}}(S(\mathbb{C}))$ is the abelian group of (global) algebraic $\mathbb{C}$-sections.
In Section 3.1 we will show that our main theorems are insensitive to finite base changes of $S$ and to isogenies. Hence, by using for instance [12, Proof of Theorem B.1 Devissages $(iv)-(vi)$] we can assume that the abelian scheme carries principal polarization and it has $-\ell$-structure.
In this paper, sometimes we also need to assume an analytic viewpoint on families of abelian varieties: we look at $S(\mathbb{C})$ and $\mathcal{A}(\mathbb{C})$ as normal quasi-projective varieties with a surjective holomorphic map $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$ whose fibers are abelian varieties, thus the zero section $\sigma_{0}:S(\mathbb{C})\to\mathcal{A}(\mathbb{C})$ and the fiberwise group operations are complex analytic. The previous family can be regarded as a geometric model of a polarized abelian variety $A$ defined over the function field $\mathbb{C}(S)$ of meromorphic functions on $S(\mathbb{C})$: from the latter perspective a rational point $\sigma$ of $A$ over $\mathbb{C}(S)$ is equivalently a complex analytic section of $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$. Moreover, we can consider normal projective compactifications $S(\mathbb{C})\hookrightarrow\overline{S}(\mathbb{C})$ and $\mathcal{A}(\mathbb{C})\hookrightarrow\overline{\mathcal{A}}(\mathbb{C})$ such that $\phi$ extends to a complex analytic map $\overline{\phi}:\overline{\mathcal{A}}(\mathbb{C})\to\overline{S}(\mathbb{C})$; suppose furthermore that the zero section extends meromorphically to $\overline{S}(\mathbb{C})$ and that the fiberwise group operations over $S(\mathbb{C})$ extends to meromorphic maps on $\overline{\mathcal{A}}(\mathbb{C})$. Then, $\overline{\phi}:\overline{\mathcal{A}}(\mathbb{C})\to\overline{S}(\mathbb{C})$ can be regarded as a geometric model of a polarized abelian variety $A$ defined over a function field $\mathbb{C}(\overline{S})$. The rational points $A(\mathbb{C}(\overline{S}))$ of $A$ over $\mathbb{C}(\overline{S})$ are then in one-to-one correspondence with the set of meromorphic sections of $\overline{\mathcal{A}}$ over $\overline{S}$ which are complex analytic over $S(\mathbb{C})$: in fact, $\sigma|_{S}:S(\mathbb{C})\to\mathcal{A}(\mathbb{C})$ is necessarily holomorphic outside a set of complex codimension $\geq 2$ on $S(\mathbb{C})$; on the other hand, over a small simply connected open set $U$ on $S(\mathbb{C})$, $\phi^{-1}(U)$ is uniformized by $U\times\mathbb{C}^{d}$. It follows by lifting $\sigma|_{U}:U\to\phi^{-1}(U)$ to $U\times\mathbb{C}^{d}$ and the Hartogs Extension Theorem for holomorphic functions that in fact $\sigma$ is everywhere holomorphic in $S(\mathbb{C})$. Notice that in particular the set of meromorphic sections of $\overline{\mathcal{A}}(\mathbb{C})$ over $\overline{S}(\mathbb{C})$ which are holomorphic over $S(\mathbb{C})$ agrees with the set of algebraic sections of $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$. By the Mordell-Weil Theorem over complex function fields, the group $A(\mathbb{C}(\overline{S}))$ is finitely generated provided that it has no constant parts.
@@
The isomorphism classes of principally polarized abelian varieties are in one-to-one correspondence with $\mathbb{H}_{g}/\textnormal{Sp}_{2g}(\mathbb{Z})$ where $\textnormal{Sp}_{2g}(\mathbb{Z})$ is a discrete subgroup of $\textnormal{Sp}_{2g}(\mathbb{R})$. To any torsion-free subgroup $\Gamma\subset\textnormal{Sp}_{2g}(\mathbb{Z})$ of finite index one can associate a universal family $\pi_{\Gamma}:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ of principally polarized abelian varieties: following Satake we call such a family $\pi_{\Gamma}$ a Kuga family of principally polarized abelian varieties. Without loss of generality, we always assume that $\Gamma$ is torsion-free. In the case of Kuga families $\pi_{\Gamma}:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ we will always take $\overline{S}_{\Gamma}(\mathbb{C})$ to be the Satake-Baily-Borel compactification: when $S_{\Gamma}(\mathbb{C})$ is a noncompact curve and the Kuga family is irreducible then $\pi_{\Gamma}:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ is isomorphic to a modular elliptic scheme, which one can compactify as an elliptic modular surface; in all other irreducible cases $\overline{S}_{\Gamma}(\mathbb{C})-S_{\Gamma}(\mathbb{C})$ are of codimension $\geq 2$ in $\overline{S}_{\Gamma}(\mathbb{C})$ so that the choice of a completion $\overline{\pi}_{\Gamma}:\overline{\mathcal{A}}_{\Gamma}(\mathbb{C})\to\overline{S}_{\Gamma}(\mathbb{C})$ is unimportant. With respect to any such compactification, the zero section $\sigma_{0}$ and the fiberwise group operations extend meromorphically. Thus, associated to each Kuga family $\pi:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ we have a well-defined modular polarized abelian variety $A_{\Gamma}$ over the modular function field $\mathbb{C}(\overline{S}_{\Gamma})$. Given a principally polarized abelian scheme $\phi:\mathcal{A}\to S$ of relative dimension $g$ with level $\ell$-structure, there is a congruence subgroup
@@
and a unique modular map $p:S\to\mathbb{H}_{g}/\Gamma=S_{\Gamma}$ such that $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$ is isomorphic to the pull-back of the Kuga family $\pi:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ by $p$. Since $\Gamma$ is torsion-free applying the extension theorem of Borel, the modular map $p:S(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ always extends to a holomorphic map $\overline{p}:\overline{S}(\mathbb{C})\to\overline{S}_{\Gamma}(\mathbb{C})$, where $\overline{S}_{\Gamma}(\mathbb{C})$ denotes the Satake-Baily-Borel compactification of $S_{\Gamma}(\mathbb{C})$; in particular, this defines a compactification $\overline{\phi}:\overline{\mathcal{A}}(\mathbb{C})\to\overline{S}(\mathbb{C})$, which is a geometric model for a polarized abelian variety $A$ over the complex function field $\mathbb{C}(\overline{S})$.
For our purposes we are interested in some vanishing results of the Mordell-Weil rank. Now, we fix once and for all a suitable level $\ell\geq 3$ such that the Kuga family $\overline{\pi}_{\Gamma}:\overline{\mathcal{A}}_{\Gamma}(\mathbb{C})\to\overline{S}_{\Gamma}(\mathbb{C})$ has no locally constant parts, i.e. there is no constant part and the same is true even when $\Gamma$ is replaced by any subgroup of finite index; in particular, observe that a base change of such a family performed via an unramified cover of the base is still a Kuga family with same properties. In [25, Main Theorem], the authors provided a finiteness result about the number of meromorphic sections of such Kuga families, which we restate in the following form:
Theorem 2.1 (Mok-To, 1993).
Let $\pi_{\Gamma}:\mathcal{A}_{\Gamma}(\mathbb{C})\rightarrow S_{\Gamma}(\mathbb{C})$ be a Kuga family of polarized abelian varieties without locally constant parts. Every algebraic section $\sigma:S_{\Gamma}(\mathbb{C})\to\mathcal{A}_{\Gamma}(\mathbb{C})$ is torsion.
Remark 2.2.
Particular instances of 2.1 are provided in [29] for special Kuga families and in [31, Theorem 5.1] in the case of elliptic modular surfaces. If $A(\mathbb{C}(\overline{S}))$ has additional structure, such as having extra endomorphisms, the classifying space $S_{\Gamma}$ should be replaced by some other modular variety; and some analogous results to 2.1 have been obtained by Silveberg in the setting of Shimura varieties (see [32]). For an alternative proof of [31, Theorem 5.1] see [9, Theorem 2.5]).
One can study the $\mathbb{Z}$-rank of $A(\mathbb{C}(\overline{S}))$ in terms of the modular map $p:S\to S_{\Gamma}$ for any Kuga family $\pi_{\Gamma}:\mathcal{A}_{\Gamma}\to S_{\Gamma}$ of polarized abelian varieties. In particular, notice that in the case when $p(S(\mathbb{C}))$ is totally geodesic and the mapping $p:S(\mathbb{C})\to p(S(\mathbb{C}))$ is unramified, then $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$ is itself a Kuga family and we still have the vanishing of the rank of the Mordell-Weil group. The general expectation is that the Mordell-Weil rank can be estimated in terms of the second fundamental form and/or the ramification locus of the modular map. In the case of a ramified covering $\overline{p}:\overline{S}\to\overline{S}_{\Gamma}$, Mok announced an upper bound of the Mordell-Weil rank in terms of the ramification divisor (see [24, Chapter 3, Theorem 2]).
@@ 2.2 Periods and abelian logarithms
Let us consider an abelian scheme $\phi:\mathcal{A}\to S$ with a zero-section as above. Each fiber $\mathcal{A}_{s}(\mathbb{C})$ is analytically isomorphic to a complex torus $\mathbb{C}^{g}/\Lambda_{s}$ and for any subset $T\subseteq S(\mathbb{C})$ we denote $\Lambda_{T}:=\bigsqcup_{s\in T}\Lambda_{s}$. The space $\operatorname{Lie}(\mathcal{A}):=\bigsqcup_{s\in S(\mathbb{C})}\operatorname{Lie}(\mathcal{A}_{s})$ has a natural structure of $g$-dimensional holomorphic vector bundle $f\colon\operatorname{Lie}(\mathcal{A})\to S(\mathbb{C})$ (it is actually a complex Lie algebra bundle). By using the fiberwise exponential maps one can define a global map $\exp\colon\operatorname{Lie}(\mathcal{A})\to\mathcal{A}$. Let $\Theta_{0}\subset\mathcal{A}$ be the image of the zero section of the abelian scheme, then obviously
@@
Clearly $S(\mathbb{C})$ can be covered by finitely many open simply connected subsets where the holomorphic vector bundle $f\colon\operatorname{Lie}(\mathcal{A})\to S(\mathbb{C})$ trivializes. Let $U\subseteq S(\mathbb{C})$ be any of such subsets and consider the induced holomorphic map $f\colon\Lambda_{U}\to U$; it is actually a fiber bundle with structure group $\textnormal{GL}(n,\mathbb{Z})$. Since $U$ is simply connected, by [11, Lemma 4.7] we conclude that $f\colon\Lambda_{U}\to U$ is a topologically trivial fiber bundle. Thus we can find $2g$ continuous sections of $f$:
@@ 2.3 Monodromy of periods
By [20, Proposition 8.1], this is the sheaf on $S$ associated to the presheaf
@@
Since $\phi:\mathcal{A}\to S$ is a smooth surjective morphism of algebraic varieties over $\mathbb{C}$, by [20, Theorem 10.4] then $\phi$ is a submersion. We can apply Ehresmann’s Lemma to conclude that the proper submersion $\phi$ is a $C^{\infty}$-fiber bundle.
@@
In the case of a universal family, a more explicit description of the monodromy action on periods is described in [16, Section 3]. Keeping the same notations as above, let us denote by $\pi:\mathfrak{A}_{g}\to\mathbb{A}_{g}$ the universal family of principally polarized $g$-dimensional abelian varieties with $-\ell$-structure for some $g\geq 1$ and $\ell\geq 3$. Let $\mathbb{H}_{g}$ denote Siegel’s upper half space, i.e., the symmetric matrices in $\textrm{Mat}_{g\times g}(\mathbb{C})$ with positive definite imaginary part. We have holomorphic uniformizing maps
@@
The previous considerations about periods of universal families allow us to obtain some results on periods of the abelian scheme $\phi:\mathcal{A}\to S$: we are able to construct period functions of $\phi:\mathcal{A}\to S$ by means of periods of the universal family $\pi:\mathfrak{A}_{g}\to\mathbb{A}_{g}$. Let $V\subseteq\mathbb{A}_{g}(\mathbb{C})$ be a simply connected open set of an open covering of $\mathbb{A}_{g}$ where holomorphic period functions do exist for the universal abelian scheme, see Equation 2; denote by $\omega^{\mathfrak{A}_{g}}_{V,1},\ldots,\omega^{\mathfrak{A}_{g}}_{V,2g}$ such local (holomorphic) period functions. Fix a base point $s\in S(\mathbb{C})$, not a ramification point for $p:S\rightarrow\mathbb{A}_{g}$, and let $x=p(s)\in\mathbb{A}_{g}(\mathbb{C})$. In a connected and simply connected neighborhood $U$ of $s$ in $S(\mathbb{C})$ such that $p(U)=V$, we can holomorphically define a basis $\omega_{U,1},\ldots,\omega_{U,2g}$ of the period lattices by the equations
@@
Recall that in our situation we have a finite map $p:S\to\mathbb{A}_{g}$ and that the monodromy group of the universal family $\textrm{Mon}(\mathfrak{A}_{g})$ is a finite index subgroup of $\textrm{Sp}_{2g}(\mathbb{Z})$. Therefore, the monodromy group $\textrm{Mon}(\mathcal{A})$ of $\mathcal{A}\to S$ is still a finite index subgroup of $\textrm{Sp}_{2g}(\mathbb{Z})$ thanks to Equation 9; in particular, $\textrm{Mon}(\mathcal{A})$ is Zariski-dense in $\textrm{Sp}_{2g}(\mathbb{Z})$.
@@
Assume that there exists a non-trivial subvector space $W$ of $\mathbb{Q}^{2g}$ which is fixed by the action of $\textrm{Sp}_{2g}(\mathbb{Z})$. Since the symplectic group $\textrm{Sp}_{2g}$ is a simple algebraic Lie group defined over $\mathbb{Q}$, by a theorem of Borel and Harish-Chandra [6, Theorem 7.8] the group $\textrm{Sp}_{2g}(\mathbb{Z})$ is a lattice in $\textrm{Sp}_{2g}(\mathbb{R})$. By Borel density theorem (see [5]) we have that $\textrm{Sp}_{2g}(\mathbb{Z})$ is Zariski dense in $\textrm{Sp}_{2g}(\mathbb{R})$, and so it is Zariski dense in $\textrm{Sp}_{2g}(\mathbb{Q})$. This implies that $W$ is fixed by the action of $\textrm{Sp}_{2g}(\mathbb{Q})$. Since the action of $\textrm{Sp}_{2g}(\mathbb{Q})$ on $\mathbb{Q}^{2g}$ is irreducible (see [19, Proposition 3.2]), we get $W=\mathbb{Q}^{2g}$ which concludes the proof.
@@
The action of the monodromy group $\textrm{Mon}(\mathcal{A})$ on the lattice of periods is irreducible.
@@
This is an easy consequence of 2.4 and 2.5.
@@
To be more precise, since the group $\textrm{Mon}(\mathcal{A})$ is non-trivial then the functions $\omega_{1},\ldots,\omega_{2g}$ cannot be all defined continuously on the whole of $S(\mathbb{C})$. Moreover, since the action of $\textrm{Mon}(\mathcal{A})$ on the lattice of periods is irreducible, neither one of $\omega_{1},\ldots,\omega_{2g}$ nor any single non-zero element of the period lattice can be defined on the whole of $S(\mathbb{C})$. This last thing is also proved in [15, Lemma 5.6] for any abelian scheme without using irriducibility properties of the monodromy action.

### v3

@@ 2.1 Setting
Let $S$ be a regular, irreducible, quasi-projective variety defined over $\mathbb{C}$; we also tacitly assume that $S$ is maximal. Let $\phi:\mathcal{A}\to S$ be an abelian scheme over $S$ of relative dimension $g\geq 1$, namely a proper smooth group $S$-scheme admitting a “zero section” $\sigma_{0}:S\to\mathcal{A}$ and whose fibers are $g$ dimensional abelian varieties. For any $s\in S(\mathbb{C})$ the fiber over $s$ is denoted by $\mathcal{A}_{s}$. We denote with $\Sigma^{\operatorname{an}}$ the sheaf over $S(\mathbb{C})$ of holomorphic sections of $\phi$ whereas $\Sigma(S)\subset\Sigma^{\operatorname{an}}(S(\mathbb{C}))$ is the abelian group of (global) algebraic $\mathbb{C}$-sections.
In Section 3.1 we will show that our main results are insensitive to finite base changes of $S$ and to isogenies. Hence, by using for instance [13, Proof of Theorem B.1 Devissages $(iv)-(vi)$] we can assume that the abelian scheme carries principal polarization and it has $-\ell$-structure.
In this paper, sometimes we also need to assume an analytic viewpoint on families of abelian varieties: we look at $S(\mathbb{C})$ and $\mathcal{A}(\mathbb{C})$ as normal quasi-projective varieties with a surjective holomorphic map $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$ whose fibers are abelian varieties, thus the zero section $\sigma_{0}:S(\mathbb{C})\to\mathcal{A}(\mathbb{C})$ and the fiberwise group operations are complex analytic. The previous family can be regarded as a geometric model of a polarized abelian variety $A$ defined over the function field $\mathbb{C}(S)$ of meromorphic functions on $S(\mathbb{C})$: from the latter perspective a rational point $\sigma$ of $A$ over $\mathbb{C}(S)$ is equivalently an algebraic section of $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$.
@@
The isomorphism classes of principally polarized abelian varieties are in one-to-one correspondence with $\mathbb{H}_{g}/\textnormal{Sp}_{2g}(\mathbb{Z})$ where $\textnormal{Sp}_{2g}(\mathbb{Z})$ is a discrete subgroup of $\textnormal{Sp}_{2g}(\mathbb{R})$. To any torsion-free subgroup $\Gamma\subset\textnormal{Sp}_{2g}(\mathbb{Z})$ of finite index one can associate a universal family $\pi_{\Gamma}:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ of principally polarized abelian varieties: following Satake we call such a family $\pi_{\Gamma}$ a Kuga family of principally polarized abelian varieties. Without loss of generality, we always assume that $\Gamma$ is torsion-free. Given a principally polarized abelian scheme $\phi:\mathcal{A}\to S$ of relative dimension $g$ with level $\ell$-structure, there is a congruence subgroup
@@
and a unique modular map $p:S\to\mathbb{H}_{g}/\Gamma=S_{\Gamma}$ such that $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$ is isomorphic to the pull-back of the Kuga family $\pi:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ by $p$. For our purposes we are interested in some vanishing results of the Mordell-Weil rank. Now, we fix once and for all a suitable level $\ell\geq 3$ such that the Kuga family $\overline{\pi}_{\Gamma}:\overline{\mathcal{A}}_{\Gamma}(\mathbb{C})\to\overline{S}_{\Gamma}(\mathbb{C})$ has no locally constant parts, i.e. there is no constant part and the same is true even when $\Gamma$ is replaced by any subgroup of finite index.
@@ 2.2 Periods and abelian logarithms
Let us consider an abelian scheme $\phi:\mathcal{A}\to S$ with a zero-section as above. Each fiber $\mathcal{A}_{s}(\mathbb{C})$ is analytically isomorphic to a complex torus $\mathbb{C}^{g}/\Lambda_{s}$ and for any subset $W\subseteq S(\mathbb{C})$ we denote $\Lambda_{W}:=\bigsqcup_{s\in W}\Lambda_{s}$. The space $\operatorname{Lie}(\mathcal{A}):=\bigsqcup_{s\in S(\mathbb{C})}\operatorname{Lie}(\mathcal{A}_{s})$ has a natural structure of $g$-dimensional holomorphic vector bundle $f\colon\operatorname{Lie}(\mathcal{A})\to S(\mathbb{C})$ (it is actually a complex Lie algebra bundle). By using the fiberwise exponential maps one can define a global map $\exp\colon\operatorname{Lie}(\mathcal{A})\to\mathcal{A}$. Let $\Theta_{0}\subset\mathcal{A}$ be the image of the zero section of the abelian scheme, then obviously
@@
Clearly $S(\mathbb{C})$ can be covered by finitely many open simply connected subsets where the holomorphic vector bundle $f\colon\operatorname{Lie}(\mathcal{A})\to S(\mathbb{C})$ trivializes. Let $U\subseteq S(\mathbb{C})$ be any of such subsets and consider the induced holomorphic map $f\colon\Lambda_{U}\to U$; it is actually a fiber bundle with structure group $\textnormal{GL}(n,\mathbb{Z})$. Since $U$ is simply connected, by [12, Lemma 4.7] we conclude that $f\colon\Lambda_{U}\to U$ is a topologically trivial fiber bundle. Thus we can find $2g$ continuous sections of $f$:
@@ 2.3 Monodromy of periods
By [21, Proposition 8.1], this is the sheaf on $S$ associated to the presheaf
@@
Since $\phi:\mathcal{A}\to S$ is a smooth surjective morphism of algebraic varieties over $\mathbb{C}$, by [21, Theorem 10.4] then $\phi$ is a submersion. We can apply Ehresmann’s Lemma to conclude that the proper submersion $\phi$ is a $C^{\infty}$-fiber bundle.
@@
In the case of a universal family, a more explicit description of the monodromy action on periods is described in [17, Section 3]. Keeping the same notations as above, let us denote by $\pi:\mathfrak{A}_{g}\to\mathbb{A}_{g}$ the universal family of principally polarized $g$-dimensional abelian varieties with $-\ell$-structure for some $g\geq 1$ and $\ell\geq 3$. Let $\mathbb{H}_{g}$ denote Siegel’s upper half space, i.e., the symmetric matrices in $\textrm{Mat}_{g\times g}(\mathbb{C})$ with positive definite imaginary part. We have holomorphic uniformizing maps
@@
The previous considerations about periods of universal families allow us to obtain some results on periods of the abelian scheme $\phi:\mathcal{A}\to S$: we are able to construct period functions of $\phi:\mathcal{A}\to S$ by means of periods of the universal family $\pi:\mathfrak{A}_{g}\to\mathbb{A}_{g}$. Let $V\subseteq\mathbb{A}_{g}(\mathbb{C})$ be a simply connected open set of an open covering of $\mathbb{A}_{g}$ where holomorphic period functions do exist for the universal abelian scheme, see Equation 3; denote by $\omega^{\mathfrak{A}_{g}}_{V,1},\ldots,\omega^{\mathfrak{A}_{g}}_{V,2g}$ such local (holomorphic) period functions. Fix a base point $s\in S(\mathbb{C})$, not a ramification point for the modular map $p:S\rightarrow\mathbb{A}_{g}$, and let $x=p(s)\in\mathbb{A}_{g}(\mathbb{C})$. In a connected and simply connected neighborhood $U$ of $s$ in $S(\mathbb{C})$ such that $p(U)=V$, we can holomorphically define a basis $\omega_{U,1},\ldots,\omega_{U,2g}$ of the period lattices by the equations
@@
Recall that the monodromy group of the universal family $\textrm{Mon}(\mathfrak{A}_{g})$ is a finite index subgroup of $\textrm{Sp}_{2g}(\mathbb{Z})$. When the modular map $p:S\to\mathbb{A}_{g}$ is finite, the monodromy group $\textrm{Mon}(\mathcal{A})$ of $\mathcal{A}\to S$ is still a finite index subgroup of $\textrm{Sp}_{2g}(\mathbb{Z})$ thanks to Equation 10; in particular, $\textrm{Mon}(\mathcal{A})$ is Zariski-dense in $\textrm{Sp}_{2g}(\mathbb{Z})$. Moreover, when the image of the modular map $p(S)$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$, the monodromy group $\textrm{Mon}(\mathcal{A})$ is still Zariski-dense in $\textrm{Sp}_{2g}(\mathbb{Z})$ by [3, Remark 10.1.2].
@@
Assume that there exists a non-trivial subvector space $W$ of $\mathbb{Q}^{2g}$ which is fixed by the action of $\textrm{Sp}_{2g}(\mathbb{Z})$. Since the symplectic group $\textrm{Sp}_{2g}$ is a simple algebraic Lie group defined over $\mathbb{Q}$, by a theorem of Borel and Harish-Chandra [7, Theorem 7.8] the group $\textrm{Sp}_{2g}(\mathbb{Z})$ is a lattice in $\textrm{Sp}_{2g}(\mathbb{R})$. By Borel density theorem (see [6]) we have that $\textrm{Sp}_{2g}(\mathbb{Z})$ is Zariski dense in $\textrm{Sp}_{2g}(\mathbb{R})$, and so it is Zariski dense in $\textrm{Sp}_{2g}(\mathbb{Q})$. This implies that $W$ is fixed by the action of $\textrm{Sp}_{2g}(\mathbb{Q})$. Since the action of $\textrm{Sp}_{2g}(\mathbb{Q})$ on $\mathbb{Q}^{2g}$ is irreducible (see [20, Proposition 3.2]), we get $W=\mathbb{Q}^{2g}$ which concludes the proof.
@@
If the image of the modular map $p(S)$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$, the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ on the lattice of periods is irreducible.
@@
This is an easy consequence of 2.2 and 2.3.
@@
To be more precise, since the group $\textrm{Mon}(\mathcal{A})$ is non-trivial then the functions $\omega_{1},\ldots,\omega_{2g}$ cannot be all defined continuously on the whole of $S(\mathbb{C})$. Moreover, when the action of $\textrm{Mon}(\mathcal{A})$ on the lattice of periods is irreducible, neither one of $\omega_{1},\ldots,\omega_{2g}$ nor any single non-zero element of the period lattice can be defined on the whole of $S(\mathbb{C})$. This last thing is also proved in [16, Lemma 5.6] for any abelian scheme over a curve without using irriducibility properties of the monodromy action.